Suppose you held a diversified portfolio consisting of a $7,500 investment in each of 20 different common stocks. The portfolio's beta is 2.25. Now suppose you decided to sell one of the stocks in your portfolio with a beta of 1.0 for $7,500 and use the proceeds to buy another stock with a beta of 2.14. What would your portfolio's new beta be? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Increase in beta as a result of net sales=(2.14-1)=1.14

Hence increase/stock=1.14/20

=0.057

Hence new beta=2.25+0.057

=2.31(Approx).